# Illuminating Darkness: Enhancing Real-world Low-light Scenes with Smartphone Images

### v1

@@ 3.1 Luminance-Chrominance Mapping
We calculated luminance ($\mathbf{I_{L}}$) and chrominance prior ($\mathbf{I_{C}}$) directly from a given low light image $\mathbf{I_{\mathbf{LL}}}\in\mathbb{R}^{H\times W\times 3}$. Hence, we extracted $\mathbf{I_{L}}$ leveraging weighted combination of color components [65, 23] and separated $\mathbf{I_{C}}$ from $\mathbf{I_{\mathbf{LL}}}$ as $\mathbf{I_{C}}=\mathbf{I_{\mathbf{LL}}}-\mathbf{I_{L}}$. Notably, directly incorporating $\mathbf{I_{L}}$ into the network enhances fine details and prevents over-brightening, while learning $\mathbf{I_{C}}$ preserves natural color fidelity and avoids desaturation. Additionally, this separation helps to target noise reduction independently in the structure (luminance) and color (chrominance) components. Therefore, we mapped $\mathbf{I_{L}}$ and $\mathbf{I_{C}}$ through mapping function $\mathbf{LC_{mapping}}$, comprising consecutive conv operation (please see 5 (b, c)) to extract features as follows:
$\mathbf{F_{L}},\mathbf{I_{B_{L}}}=\mathbf{LC_{mapping}}(\mathbf{I_{\mathbf{LL}}},\mathbf{I_{L}})$ (2)
where $\mathbf{F_{L}}$, $\mathbf{I_{B_{L}}}\in\mathbb{R}^{H\times W\times 40}$ are luminance feature and luminance boosted image (i.e., lit-up image [4]). Similarly, we obtain chrominance feature $\mathbf{F_{C}}$ and chrominance boosted image $\mathbf{I_{BC}}$ stated in the above en. 2. Notably, the previous SLLIE method [4] based on Retinex theory calculated boosted image $\mathbf{I_{B}}$ with illumination features. We found that such direct manipulation utilizing illumination prior pushes the deep networks to focus on the luminance information, ignoring the important chrominance attributes and being prone to over-enhancement.
Figure 5: Overview of proposed TFFormer. (a) Overall architecture, (b) Chrominance mapping, (c) Luminance mapping, (d) LCCAB, and (e) LCGRB.
@@ 3.2 LC Encoder-Decoder
The proposed TFFormer encodes LC attributes using the dedicated encoder. The luminance encoder learns to map $\{\mathbf{I_{BL}},\mathbf{F_{L}}\}$ as $\{\mathbf{I^{\ast}_{BL}},\mathbf{F^{\ast}_{L}}\}\ =\mathbf{E_{L}}\in\mathbb{R}^{\left(\frac{H}{8}\right)\times\left(\frac{W}{8}\right)\times 320}$. Similarly, we obtain chrominance encoder features as $\{\mathbf{I^{\ast}_{BC}},\mathbf{F^{\ast}_{C}}\}=\mathbf{E_{C}}\in\mathbb{R}^{\left(\frac{H}{8}\right)\times\left(\frac{W}{8}\right)\times 320}$. We refined the encoded LC features further, leveraging our proposed cross-attention module (sec. 3.3) for inter-attribute correlation to generate $\mathbf{A_{LC}}$. We combine correlated feature ($\mathbf{A_{LC}}$) along with $\mathbf{E_{L}}$ and $\mathbf{E_{C}}$ to reconstruct an intermediate image $\mathbf{I_{rec}}\in\mathbb{R}^{H\times W\times 3}$. To leverage LC properties and salient feature extraction effectively, we introduce LCGAB (an extension of IGAB [4]) in the encoder-decoder blocks (see Fig. 5).
@@ 3.3 LC Cross Attention Block (LCCAB)
We leverage LCCAB to refine and correlate luminance and chrominance features. Therefore, we leverage multihead self-attention (S) [4, 52, 48, 45] to refine encoded $(\mathbf{E_{L}})$ and $(\mathbf{E_{C}})$ features such as $\mathbf{A_{L}}=S(\mathbf{E_{L}})$ and $\mathbf{A_{C}}=S(\mathbf{E_{C}})$. Here $(\mathbf{A_{L}},\mathbf{A_{C}})\in\mathbb{R}^{HW\times 320}$. We further refined and correlated the luminance attention $\mathbf{A_{L}}$ and chrominance attention $\mathbf{A_{C}}$ with multi-head cross-attention (CA) block to enhance inter-domain correlation and capture global attention over LC encoding as:
$\mathbf{A_{LC}}=\operatorname{CA}(\mathbf{A_{L}},\mathbf{A_{C}})=\operatorname{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (3)
Here $(\mathbf{A_{L}},\mathbf{A_{C}})\in\mathbb{R}^{HW\times C}$ is further split into k heads such that $(\mathbf{A_{L}},\mathbf{A_{C}})\in\mathbb{R}^{HW\times\mathbf{d_{k}}},\mathbf{d_{k}}=\frac{C}{k}$ where k=8. $\mathbf{Q},\mathbf{K},\mathbf{V}$ are projections of $\mathbf{A_{L}}$ and $\mathbf{A_{C}}$ into query, key, and value spaces. Here, LC cross-attention captures long-range dependencies between the luminance and chrominance domains, improving global illumination and color constancy during the reconstruction of $\mathbf{I_{rec}}$.
@@ 3.4 LC Guided Reconstruction Block (LCGRB)
The proposed LCGRB refine the intermediate image as $\mathrm{R}:\mathbf{I_{\mathbf{L_{Rec}}}\to\mathbf{I_{Ref}}}$. Therefore, it extract the reconstructed luminance ($\mathbf{I^{\ast}_{L}}$) and reconstructed chrominance ($\mathbf{I^{\ast}_{C}}$) (similar to Sec. 3.1) and map them as $\mathbf{I^{\ast}_{L}}\in\mathbb{R}^{H\times W\times 20}$ and $\mathbf{I^{\ast}_{C}}\in\mathbb{R}^{H\times W\times 20}$. We leverage $\textit{conv}_{3\times 3}$ for high-dimensional projections of LC properties to map combined feature $\mathbf{R_{C}}\in\mathbb{R}^{H\times W\times 40}$. We feed $\mathbf{R_{C}}$ and decoded feature $\mathbb{F_{D}}^{H\times W\times 40}$ to LCGAB to obtain the refined output ($\mathbf{I_{ref}}$), as shown in Fig. 5 (d). LCGRB helps TFFormer to refine its reconstructions globally by leveraging LC properties without introducing any further training phase.
3.5 LC Perceptual Loss (LCPL)
The core idea of TFFormer is to leverage LC attributes within a deep network. To enhance alignment and refinement, we compute per-pixel LC distance by comparing ground truth and generated LC components. Our proposed LCPL enforces consistency between intermediate and refined representations, improving color fidelity and structural details. The per-pixel LC distance is defined as:
$\mathcal{D}_{LC}=\frac{1}{N}\sum_{i=1}^{N}\left(\|\mathbf{I_{G,L}^{(i)}}-\mathbf{I_{rec,L}^{(i)}}\|_{1}+\|\mathbf{I_{G,C}^{(i)}}-\mathbf{I_{rec,C}^{(i)}}\|_{1}\right)$ (4)
where $\mathbf{I_{G,L}}$ and $\mathbf{I_{rec,L}}$ represent the luminance components of the ground truth and reconstructed images, respectively. $\mathbf{I_{G,C}}$ and $\mathbf{I_{rec,C}}$ represent the chrominance components. $N$ is the total number of pixels. The LC loss is formulated to guide the network in learning LC representations effectively:
$\mathcal{L}_{LC}=\mathcal{D}_{LC}(\mathbf{I_{G},I_{rec}})+\mathcal{D}_{LC}(\mathbf{I_{G},I_{ref}})$ (5)
We leverage $I_{ref}$ and $I_{rec}$ images to calculate our perceptual loss, enhance gradient flow, stabilize training, and enable progressive refinement for improved performance.
@@ 4 Experiments
Low-light Extreme Low-light LSD
Model Type Indoor Outdoor Indoor Outdoor Average
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Retinexnet [49] 14.57 0.6327 0.2611 13.70 0.5489 0.3150 15.63 0.5754 0.3064 14.34 0.4871 0.3875 14.56 0.5610 0.3175
Kind [61] 13.55 0.6717 0.2665 13.47 0.6287 0.3174 14.21 0.6320 0.3001 13.25 0.6047 0.3521 13.62 0.6343 0.3090
Kind+ [62] 13.49 0.6630 0.2871 13.17 0.6138 0.3411 13.94 0.6141 0.3275 13.13 0.5899 0.3733 13.43 0.6202 0.3322
DeepUPE [47] 15.73 0.5874 0.3008 15.03 0.5620 0.2841 14.99 0.4612 0.4104 16.30 0.4378 0.4131 15.51 0.5121 0.3521
SNRNet [52] DLI 15.14 0.6326 0.3776 14.82 0.3510 0.7161 13.05 0.5377 0.4450 15.43 0.4646 0.5720 14.61 0.4965 0.5277
MIRNet [57] 16.15 0.7206 0.2284 14.01 0.5600 0.3632 15.51 0.6360 0.2700 17.20 0.5811 0.3085 15.72 0.6244 0.2925
Uformer [48] 13.04 0.6578 0.2672 12.40 0.5423 0.3635 11.93 0.5575 0.3092 13.46 0.5006 0.3593 12.71 0.5646 0.3248
IAT [7] 13.34 0.6557 0.2819 13.31 0.4680 0.4320 14.82 0.4978 0.3944 13.07 0.5600 0.3583 13.64 0.5454 0.3667
Retinexformer [4] 17.77 0.7338 0.2138 17.40 0.5828 0.3199 17.53 0.6591 0.2545 17.40 0.5828 0.3199 17.53 0.6396 0.2770
TFFormer 19.99 0.7620 0.2129 17.72 0.6830 0.2557 18.86 0.7121 0.2430 19.95 0.6555 0.3124 19.13 0.7031 0.2560
Retinexnet [49] 14.11 0.5456 0.3465 13.21 0.5234 0.3710 14.44 0.4252 0.5545 13.50 0.4078 0.5077 13.81 0.4755 0.4449
Kind [61] 13.49 0.5330 0.3495 13.09 0.5701 0.3280 13.43 0.4331 0.4819 13.13 0.5532 0.3699 13.28 0.5223 0.3823
Kind+ [62] 13.22 0.5283 0.3662 12.70 0.5667 0.3405 13.62 0.4214 0.4949 12.90 0.4476 0.4625 13.11 0.4910 0.4160
DeepUPE [47] 14.11 0.3529 0.5667 13.42 0.4061 0.4229 12.79 0.2125 0.7857 14.01 0.2799 0.6351 13.58 0.3128 0.6026
SNRNet [52] NLI 16.66 0.5793 0.4208 14.95 0.3025 0.7497 14.04 0.4882 0.5207 16.60 0.4110 0.6183 15.56 0.4453 0.5774
MIRNet [57] 14.82 0.6551 0.2523 13.64 0.5297 0.3274 12.80 0.5291 0.3615 16.74 0.5524 0.3300 14.50 0.5666 0.3178
Uformer [48] 13.85 0.6570 0.2628 13.07 0.5654 0.3315 12.53 0.5591 0.3488 14.13 0.5325 0.3609 13.39 0.5785 0.3260
IAT [7] 15.35 0.6506 0.2612 13.92 0.5248 0.3434 13.33 0.5308 0.3980 15.88 0.5062 0.3671 14.62 0.5531 0.3424
Retinexformer [4] 18.00 0.6934 0.2470 14.99 0.5531 0.3534 17.47 0.6094 0.3387 17.89 0.5822 0.3564 17.08 0.6095 0.3239
TFFormer 19.34 0.7198 0.2445 17.91 0.6837 0.2394 19.27 0.6923 0.3048 20.34 0.6401 0.3034 19.21 0.6840 0.2730
Table 2: Quantitative comparison of SLLIE methods on proposed LSD. Our proposed TFFormer outperforms the existing methods by a notable margin. The best and the second-best results are highlighted in red and blue, respectively. Best viewed in zoom.

### v2

@@ 3.1 Luminance-Chrominance Mapping
Given a low-light RGB image $\mathbf{I}_{LL}\in\mathbb{R}^{H\times W\times 3}$, we separate it into a luminance map $\mathbf{I}_{L}$ and chrominance prior $\mathbf{I}_{C}$. The luminance is a weighted sum of the color channels following [58, 20], and the chrominance is obtained by subtracting: $\mathbf{I}_{C}=\mathbf{I}_{LL}-\mathbf{I}_{L}$. Injecting raw $\mathbf{I}_{L}$ preserves structural details and curbs over-brightening, while learning on $\mathbf{I}_{C}$ safeguards color fidelity and prevents desaturation. This separation enables independent noise suppression in structure-sensitive (luminance) and color-sensitive (chrominance) domains. Each component passes through an identical mapping block, $\mathrm{LC}{\text{map}}(\cdot)$, built from three conv–BN–GELU layers (Fig. 5 (a,b)). The module outputs a feature tensor and ”boosted” image:
$(\mathbf{F}_{L},\mathbf{I}_{B_{L}})$ $=\mathrm{LC}_{\text{map}}(\mathbf{I}_{LL},\mathbf{I}_{L})$ (2)
$(\mathbf{F}_{C},\mathbf{I}_{B_{C}})$ $=\mathrm{LC}_{\text{map}}(\mathbf{I}_{LL},\mathbf{I}_{C})$
where $\mathbf{F}_{L}$, $\mathbf{F}_{C}$, $\mathbf{I}_{B_{L}}$, $\mathbf{I}_{B_{C}}$in $\mathbb{R}^{H\times W\times 40}$. Retinex-based SLLIE models such as Retinexformer [3] compute boosted images directly from illumination features, biasing toward luminance and yielding colour over-enhancement. Our decoupled design retains chrominance information, leading to balanced enhancement with fewer colour artefacts.
Figure 5: Overview of the proposed TFFormer architecture. (a), (b) LC mapping modules extract and boost luminance and chrominance features. (c), (d) Dedicated encoders process the respective LC features. (e) LC Cross-Attention Block (LCCAB) fuses luminance and chrominance representations. (f) A shared decoder reconstructs an intermediate output. (g) The LC Guided Refinement Block (LCGRB) further enhances the image using LC-aware attention.
@@ 3.2 LC Encoder-Decoder
TFFormer adopts an LC Encoder-Decoder architecture, where luminance and chrominance features are encoded separately using modality-specific encoders and fused during decoding (Fig. 5(c,d)). The luminance and chrominance encoders, $\mathcal{E}_{L}(\cdot)$ and $\mathcal{E}_{C}(\cdot)$, comprise multiple LC Guided Attention Blocks (LCGABs), which extract structure and color-sensitive features guided by semantic cues. Given boosted inputs $(I_{B},F)\in\mathbb{R}^{HW\times C}$ from the LC mapping module, each LCGAB splits $I_{B}$ into $k$ attention heads, $I_{B}^{i}\in\mathbb{R}^{HW\times d_{k}}$, where $d_{k}=C/k$. Each head computes query, key, and value matrices:
$Q_{i}=I_{B}^{i}W_{i}^{Q},\quad K_{i}=I_{B}^{i}W_{i}^{K},\quad V_{i}=I_{B}^{i}W_{i}^{V},$ (3)
where $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}\in\mathbb{R}^{d_{k}\times d_{k}}$ are head-specific parameters. The scaled dot-product attention is computed using a learnable scalar temperature $\alpha_{i}$ [3]:
$\mathrm{Attn}(I_{B},F)=\operatorname{Softmax}\left(\tfrac{Q_{i}K_{i}^{\top}}{\alpha_{i}}\right)V_{i}\odot F.$ (4)
Here, $\odot$ denotes element-wise multiplication, where the guidance LC feature $F$ modulates the attended output. LC encoders produce enhanced feature sets:
$E_{L}$ $=\mathcal{E}_{L}(I_{B_{L}},F_{L})=(I_{B_{L}},F_{L}),$ (5)
$E_{C}$ $=\mathcal{E}_{C}(I_{B_{C}},F_{C})=(I_{B_{C}},F_{C}),$
where $E_{L},E_{C}\in\mathbb{R}^{\tfrac{H}{8}\times\tfrac{W}{8}\times 320}$ represent the refined luminance and chrominance features. A cross-attention module (Sec. 3.3) merges $E_{L}$ and $E_{C}$ into a fused representation $A_{LC}$. The shared decoder upsamples $A_{LC}$ through a symmetric stack of LCGABs, integrating skip connections from $E_{L}$ and $E_{C}$ to preserve fine details. The decoder outputs the intermediate enhanced RGB image $I_{\text{rec}}\in\mathbb{R}^{H\times W\times 3}$ with adaptively fused luminance and chrominance.
$I_{\text{rec}}=\mathcal{D}(A_{LC},E_{L},E_{C}),$ (6)
where $\mathcal{D}(\cdot)$ denotes the joint decoder that performs attention-guided upsampling and fusion using the encoder outputs.
@@ 3.3 LC Cross Attention Block (LCCAB)
Separate luminance and chrominance encoding helps preserve structural details and enables domain-specific denoising. However, this can hinder cross-domain interaction, leading to spatial and chromatic inconsistencies. To address this, we introduce the LCCAB, which adaptively fuses the LC features for perceptual coherence. As shown in Fig. 5 (e), LCCAB applies multi-head self-attention to luminance and chrominance tokens independently, producing $\mathbf{A_{L}}=\text{Attn}(\mathbf{E_{L}})$ and $\mathbf{A_{C}}=\text{Attn}(\mathbf{E_{C}})$. Cross-attention is then performed with $\mathbf{A_{L}}$ as queries and $\mathbf{A_{C}}$ as keys and values, resulting in the fused representation $\mathbf{A_{LC}}=\text{CA}(\mathbf{A_{L}},\mathbf{A_{C}})$. This enhances inter-domain dependencies, improving illumination and chroma alignment in the reconstructed image $\mathbf{I_{rec}}$.
@@ 3.4 LC Guided Reconstruction Block (LCGRB)
The LCGRB refines the intermediate image $\mathbf{I_{\mathbf{L_{Rec}}}}\to\mathbf{I_{Ref}}$ by extracting reconstructed luminance ($\mathbf{I^{\ast}_{L}}$) and chrominance ($\mathbf{I^{\ast}_{C}}$) components (Section 3.1), mapped to $\mathbf{I^{\ast}_{L}},\mathbf{I^{\ast}_{C}}\in\mathbb{R}^{H\times W\times 20}$ using $3\times 3$ convolution refinement. They are concatenated to form $\mathbf{R_{LC}}\in\mathbb{R}^{H\times W\times 40}$. The LCGAB processes $\mathbf{R_{LC}}$ to produce the refined output $\mathbf{I_{ref}}$ (Fig. 5(e)). This enables global reconstruction refinement by leveraging LC properties without additional training phases. Ablation studies in Section 4.4 validate the effectiveness of all proposed blocks.
3.5 LC Consistency Guidance (LCCG)
TFFormer exploits LC characteristics by computing a per-pixel $\ell_{1}$ distance between the predicted and ground-truth LC components to mitigate over-smoothing [32]. This distance is applied to both the intermediate reconstruction $I_{\text{rec}}$ and the final refinement $I_{\text{ref}}$. LCCG enforces consistency between the two stages, enhances color fidelity, sharpens structural details, and provides smoother gradients for stable training and progressive refinement.
$\mathcal{D}_{\text{LC}}\;=\;\frac{1}{N}\sum_{i=1}^{N}\Bigl{(}\lVert I^{\text{gt}}_{L,(i)}-I^{\text{rec}}_{L,(i)}\rVert_{1}\;+\;\lVert I^{\text{gt}}_{C,(i)}-I^{\text{rec}}_{C,(i)}\rVert_{1}\Bigr{)}$
where $I^{\text{gt}}_{L}$ and $I^{\text{rec}}_{L}$ (respectively $I^{\text{gt}}_{C}$ and $I^{\text{rec}}_{C}$) denote the luminance (respectively chrominance) channels of the ground-truth and reconstructed images, and $N$ is the number of pixels. We reuse the same distance for the refinement output, giving the overall loss
$\mathcal{L}_{\text{LC}}=\mathcal{D}_{\text{LC}}\!\bigl{(}I^{\text{gt}},\,I^{\text{rec}}\bigr{)}+\mathcal{D}_{\text{LC}}\!\bigl{(}I^{\text{gt}},\,I_{\text{ref}}\bigr{)}$ (8)
Low-light Extreme Low-light LSD
Compexity Indoor Outdoor Indoor Outdoor Average
Model Scene Flops (G) Param. (M) PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Retinexnet [42] 136.00 0.84 14.57 0.6327 0.2611 13.70 0.5489 0.3150 15.63 0.5754 0.3064 14.34 0.4871 0.3875 14.56 0.5610 0.3175
Kind [54] 64.73 8.01 13.55 0.6717 0.2665 13.47 0.6287 0.3174 14.21 0.6320 0.3001 13.25 0.6047 0.3521 13.62 0.6343 0.3090
Kind+ [55] 59.41 8.27 13.49 0.6630 0.2871 13.17 0.6138 0.3411 13.94 0.6141 0.3275 13.13 0.5899 0.3733 13.43 0.6202 0.3322
DeepUPE [40] 0.04 0.59 15.73 0.5874 0.3008 15.03 0.5620 0.2841 14.99 0.4612 0.4104 16.30 0.4378 0.4131 15.51 0.5121 0.3521
SNRNet [45] 23.96 39.12 15.14 0.6326 0.3776 14.82 0.3510 0.7161 13.05 0.5377 0.4450 15.43 0.4646 0.5720 14.61 0.4965 0.5277
MIRNet [51] 816.00 31.79 16.15 0.7206 0.2284 14.01 0.5600 0.3632 15.51 0.6360 0.2700 17.20 0.5811 0.3085 15.72 0.6244 0.2925
Uformer [41] 12.00 5.29 13.04 0.6578 0.2672 12.40 0.5423 0.3635 11.93 0.5575 0.3092 13.46 0.5006 0.3593 12.71 0.5646 0.3248
IAT [6] 1.44 0.09 13.34 0.6557 0.2819 13.31 0.4680 0.4320 14.82 0.4978 0.3944 13.07 0.5600 0.3583 13.64 0.5454 0.3667
Retinexformer [3] 39.16 3.38 16.83 0.7238 0.2286 15.76 0.5900 0.3551 17.97 0.6746 0.2495 17.73 0.5945 0.3123 17.07 0.6457 0.2864
GSAD [13] 67.02 17.17 18.31 0.7307 0.2195 16.55 0.5934 0.2916 18.52 0.7042 0.2425 17.06 0.6232 0.2993 17.61 0.6629 0.2632
Diff-Retinex [50] 492.47 80.70 14.85 0.6235 0.2795 14.89 0.5229 0.3663 16.16 0.5993 0.3302 16.35 0.5070 0.4143 15.56 0.5632 0.3476
LYT [2] 1.70 0.04 17.26 0.7103 0.2122 16.41 0.5533 0.3776 17.31 0.6386 0.2789 18.41 0.5725 0.3363 17.35 0.6187 0.3012
HVI [47] 8.13 1.97 17.30 0.7200 0.2325 16.68 0.5690 0.3006 16.63 0.6692 0.2860 17.63 0.6009 0.3325 17.06 0.6398 0.2879
TFFormer DLL 34.59 5.87 19.28 0.7766 0.2083 17.71 0.6576 0.3061 19.82 0.7363 0.2335 19.70 0.6660 0.2954 19.13 0.7091 0.2609
Retinexnet [42] 136.00 0.84 14.11 0.5456 0.3465 13.21 0.5234 0.3710 14.44 0.4252 0.5545 13.50 0.4078 0.5077 13.81 0.4755 0.4449
Kind [54] 64.73 8.01 13.49 0.5330 0.3495 13.09 0.5701 0.3280 13.43 0.4331 0.4819 13.13 0.5532 0.3699 13.28 0.5223 0.3823
Kind+ [55] 59.41 8.27 13.22 0.5283 0.3662 12.70 0.5667 0.3405 13.62 0.4214 0.4949 12.90 0.4476 0.4625 13.11 0.4910 0.4160
DeepUPE [40] 0.04 0.59 14.11 0.3529 0.5667 13.42 0.4061 0.4229 12.79 0.2125 0.7857 14.01 0.2799 0.6351 13.58 0.3128 0.6026
SNRNet [45] 23.96 39.12 16.66 0.5793 0.4208 14.95 0.3025 0.7497 14.04 0.4882 0.5207 16.60 0.4110 0.6183 15.56 0.4453 0.5774
MIRNet [51] 816.00 31.79 14.82 0.6551 0.2523 13.64 0.5297 0.3274 12.80 0.5291 0.3615 16.74 0.5524 0.3300 14.50 0.5666 0.3178
Uformer [41] 12.00 5.29 13.85 0.6570 0.2628 13.07 0.5654 0.3315 12.53 0.5591 0.3488 14.13 0.5325 0.3609 13.39 0.5785 0.3260
IAT [6] 1.44 0.09 15.35 0.6506 0.2612 13.92 0.5248 0.3434 13.33 0.5308 0.3980 15.88 0.5062 0.3671 14.62 0.5531 0.3424
Retinexformer [3] 39.16 3.38 17.16 0.6903 0.2493 15.14 0.5908 0.3297 17.22 0.6216 0.3140 17.40 0.5585 0.3549 16.73 0.6153 0.3120
GSAD [13] 67.02 17.17 17.06 0.6232 0.2993 16.21 0.5806 0.2929 17.89 0.6601 0.2907 16.45 0.5723 0.3589 16.90 0.6091 0.3104
Diff-Retinex [50] 492.47 80.70 14.73 0.4990 0.3743 14.87 0.4745 0.3342 15.13 0.3905 0.5243 16.05 0.4131 0.4494 15.20 0.4443 0.4206
LYT [2] 1.70 0.04 16.33 0.5886 0.2926 15.72 0.5070 0.3837 16.27 0.4915 0.4289 16.71 0.4507 0.4172 16.26 0.5095 0.3806
HVI [47] 8.13 1.97 16.43 0.6593 0.2816 15.77 0.5455 0.3161 16.55 0.6384 0.3360 16.42 0.5443 0.3842 16.29 0.5969 0.3295
TFFormer NLL 34.59 5.87 18.99 0.7470 0.2231 17.28 0.6799 0.2823 18.94 0.7113 0.2571 19.75 0.6598 0.3018 18.74 0.6995 0.2661
Table 2: Quantitative Benchmark on LSD Dataset. Our proposed TFFormer outperforms the existing methods by a notable margin. The best and the second-best results are highlighted in red and blue, respectively. Best viewed in zoom.